OpenStudy (anonymous):

If 15.00g of Mg is reacted with an excess of H2SO4, MgSO4 and H2 are produced. The theoretical yield of MgSO4 is 75g. A student performed the experiment and weighed out 71g. What is the percent error? A.1% B 4% C. 5% D. 95% I'm not looking for a direct answer, just help finding it.

OpenStudy (anonymous):

You don't need to do any stoichiometry here. \[\color{green} { \begin{align*}\text{% error } &= \frac{\mid \text{theoretical yield } - \text{actual yield}\mid}{\text{theoretical yield}} \times 100 \\ &= \frac{75 - 71}{75} \times 100 \end{align*}} \]